# Elastic wavepackets crossing a space-time interface

### v1

@@ Abstract
The interaction between waves and evolving media challenges traditional conservation laws. We experimentally investigate the behavior of elastic wave packets crossing a moving interface that separates two media with distinct propagation properties, observing the non-conservation of wavelength and frequency. Our experimental setup employs an elastic strip whose local stretching can be dynamically altered by pulling one end at a constant velocity. By demonstrating that this dynamic configuration creates a spatio-temporal interface traveling along the strip, we confirm theoretical predictions regarding observed shifts when a wave packet crosses this interface.
@@
Despite the extensive theoretical explanations and studies of the underlying physics [26, 27, 28, 29, 30, 31, 32], these phenomena lack direct experimental evidence. In this article, we present an easy-to-implement experiment that allows the observation of a unique event on a macroscopic scale: the crossing of a spatio-temporal interface by a wave. For this, we start with a soft elastic strip and suddenly displace one of its two ends. This creates a moving front of transition between a region at rest and an elongated one. Then, we study the passage of a flexural wave packet through this interface in two possible configurations: the wave and the interface propagate either in opposite directions or in the same direction. Our experimental results, which are supported with a theoretical model and demonstrate that both frequency and momentum are shifted by this interaction.
@@
Figure 1: Experimental setup and media characterization — A- An elastic strip is attached at the upper end and pulled at a constant velocity $V_{0}$ at the lower end. The experiment is recorded with a camera at a frame rate of 1 kHz. The transverse displacement is extracted by tracking the position of the beam’s edge, and the axial deformation by tracking the position in $x_{1}$ of darker spots on the edge. B- Spatio-temporal representation of the strip elongation. The bottom edge starts moving a $t=0$ with a speed $V_{0}$, and an interface propagates leftward at speed $V_{\textrm{interface}}$ (as sketched below). The darker spots are shown as red circles and their position evolution are shown. The red arrow indicates the spatio-temporal interface moving at $V_{\textrm{interface}}$, with blue arrows highlighting the medium’s stretching and drift relative to the laboratory frame. C Map of the extracted stretch ratio $\lambda$. D- Repeating the experiment for different $V_{0}$, a linear relation gives $V_{\textrm{interface}}=8.2$ m/s. E Map of the extracted particle velocity.
@@
$\frac{EI}{A}\frac{\partial^{4}u_{3}}{\partial{x_{1}}^{4}}-\sigma\frac{\partial^{2}u_{3}}{\partial{x_{1}}^{2}}-\rho\frac{\partial^{2}u_{3}}{\partial t^{2}}=0,$ (1)
with $u_{3}$ denoting the out-of plane displacement of the beam, $E$ the Young modulus of the material, $A=wh$ the beam’s cross-section, $I=wh^{3}/12$ the second moment of area, $\sigma$ the applied longitudinal stress (or tension) and $\rho$ the mass density of the material. This equation immediately reveals that $\sigma$ acts as a tuning parameter: as $\sigma$ varies, a competition between the first and second terms explicitly emerges. At low stress, the propagation is governed by the bending moment $EI/A$. But, as the stress increases the tension term begins to dominate [34]. Alternatively to the stress, this criterion can also be written in terms of elongational strain $\lambda$, invoking Hooke’s law $\sigma=E(\lambda-1)$.
@@
Figure 2: Crossing a space-time interface — Space-time diagram (and reciprocal representation) of a flexural wave packet crossing a space-time interface propagating in the opposite direction (A) or in the same direction (B). The strip is pulled at 4 m/s in (A) and 3.3 m/s in (B). The non-conservation of both wavelength (or wavenumber) and period (or frequency) for the transmitted waves is demonstrated. Below are represented the spatio-temporal Fourier transforms of the same data. Superimposed are shown the theoretical dispersion curves before (blue) and after the interface (green).
@@
Figure 1.C shows the strip’s elongation map as a function of space and time. It demonstrates a spatiotemporal transition from initial rest to a stretched configuration. Because $\lambda$ explicitly depends on $V_{0}$, the contrast in elongation between the two regions can be controlled by tuning the motor velocity $V_{0}$. And, as discussed earlier, the amount of elongation directly sets the velocity of the in-plane flexural waves. In comparison to the theoretical model, the interface is not as sharp as a step function but rather exhibits a slower variation from 1 to 1.16. It can be attributed to the acceleration of the motor and to the rheology of the material [39, 40, 34]. A systematic extraction of the stretch ratio for various values of $V_{0}$ is presented in Figure 1.D. Not only it confirms the possibility to create a wide variety of elongation ratios, but the agreement with equation (4) allows to retrieve a precise value of $V_{\textrm{interface}}=8.2$ m/s.
A second crucial aspect to this experimental setup involves a peculiarity of the stretched medium. According to equation (3), after the interface passage, the medium experiences a uniform motion at velocity $V_{0}$. Consequently, not only the medium is stretched but it also undergoes a uniform translation. This appears readily on Figure 1.B, where red dots and blue arrows indicate linear translation to the right. The experimental drifting velocity obtained from the experiment is color-coded in Figure 1.E. After the interface passage, a homogeneous medium drifting at the constant velocity $V_{0}$ is observed in the upper right corner of the figure. This translation significantly affects the propagation of the transverse wave when observed in the laboratory frame. Specifically, due to the translation of the wave’s support, a Doppler shift needs to be accounted for:
@@
These results exemplify the paradigm shift introduced by considering a moving interface. Generally, the conservation of frequency (or equivalently, the temporal period) when crossing a spatial interface is justified by invoking the invariance by time translation along with the linearity of the medium. However, recent experiments involving a temporal interface [44, 45], echoing previous theoretical developments [46, 47], have revealed that the presence of a reflected time-reversed wave can be attributed to the conservation of momentum (or wavelength). In our case, none of these quantities is conserved. It underscores the fundamental changes brought about by the moving interface [30]. This phenomenon can be explained by the following general conservation law:
@@
where $\omega$ corresponds to the angular frequency and $k$ corresponds to the momentum of the wave. It is worth noting that negative wavenumbers are required when the wave propagates in the negative direction.
Despite the broken time and space invariance in this system, the 2D Fourier Transform still yields valuable information about the spectral contents. The results are depicted below the spatio-temporal maps of the waves in Figure 2. For each scenario, two peaks in the plane reciprocal plane are observed. For the wave propagating in the negative direction, the wavenumbers naturally correspond to negative values. For the strip at rest (blue) as well as for the stretched and moving one (green), all the peaks coincide with the full 3D calculation of the dispersion relations [34, 48]. The transition in the dispersion relation diagram between Medium 1 and Medium 2 also demonstrates these momentum and frequency shifts, aligning with equation (6) along the line of slope $-V_{\textrm{interface}}$.
Eventually, the experiment is repeated for various pulling velocities and wavepacket’s frequencies. For each experiment 2D space-time Fourier transform are applied and the relevant spots of the Fourier transforms are kept. All the results are accumulated and are summarized on the same axes in Figure 3. The initial wavepackets are all aligned with the dispersion relation of flexural waves in a strip almost at rest (black line). A slight difference between the negative and positive wavenumbers is visible because of the gravity that exerts a smaller elongation of the strip at its bottom compared to the top. After crossing the interface, all wavepackets show a change in frequency and wavenumbers. The transmitted wave ends on the dispersion curve of the medium 2 which depends on the pulling velocity $V_{0}$ with a good agreement with the theoretical predictions (lines on the figure). The transition between the incident wave packet and the transmitted one follows the equation (6) as indicated by the red arrows that are all parallel with the same slope of $-V_{\textrm{interface}}$.
@@
This simple macroscopic experiment effectively demonstrates the key feature of a spatio-temporal interface: the non-conservation of momentum and frequency. However, some comments regarding this platform are needed.
@@
Secondly, the role of advection in this experiment contributes to the significant frequency shift observed in scenario A. This property can be found in other wave phenomena, such as gravity-capillary waves at the surface of moving fluids [50, 51], acoustic waves in flowing pipes [41, 43], or elastic waves in waveguides coupled to flowing fluids like arteries [52]. These examples could also be promising candidates for investigating static boundaries in the laboratory frame, which would be in translation in the moving frame.
@@
This article presents an experiment where a wave packet crosses a moving interface separating two media with distinct propagation properties. This experiment relies on the propagation of a transverse wave in an elastic strip which is dynamically stretched. The observed transmitted wave in our setup exhibits both a frequency shift and a momentum shift, driven by the velocity of the interface and the propagation characteristics of the two media.
@@
This work has received support under the program “Investissements d’Avenir” launched by the French Government and partially by the Simons Foundation/Collaboration on Symmetry-Driven Extreme Wave Phenomena.

### v2

@@ Abstract
The interaction between waves and evolving media challenges traditional conservation laws. We experimentally investigate the behavior of elastic wave packets crossing a moving interface that separates two media with distinct propagation properties, observing the non-invariance of wavelength and frequency. Our experimental setup employs an elastic strip whose local stretching can be dynamically altered by pulling one end at a constant velocity. By demonstrating that this dynamic configuration creates a spatio-temporal interface traveling along the strip, we confirm theoretical predictions regarding observed shifts when a wave packet crosses this interface.
@@
Despite the extensive theoretical explanations and studies of the underlying physics [26, 27, 28, 29, 30, 31, 32], these phenomena lack direct experimental evidence. In this article, we present an easy-to-implement experiment that allows the observation of a unique event on a macroscopic scale: the crossing of a spatio-temporal interface by a wave. For this, we start with a soft elastic strip and suddenly displace one of its two ends. This creates a moving front of transition between a region at rest and an elongated one. Then, we study the passage of a flexural wave packet through this interface in two possible configurations: the wave and the interface propagate either in opposite directions or in the same direction. Our experimental results, which are supported with a theoretical model and demonstrate that both frequency and wavenumber are shifted by this interaction.
@@
Figure 1: Experimental setup and media characterization — A- An elastic strip is attached at the upper end and pulled at a constant velocity $V_{0}$ at the lower end. The experiment is recorded with a camera at a frame rate of 1 kHz. The transverse displacement is extracted by tracking the position of the beam’s edge, and the axial deformation by tracking the position in $x_{1}$ of darker spots on the edge. B- Spatio-temporal representation of the strip elongation. The bottom edge starts moving a $t=0$ with a speed $V_{0}$ (here 1.4 m/s), and an interface propagates leftward at speed $V_{\textrm{interface}}$ (as sketched below). The darker spots are shown as red circles and their position evolution are shown. The red arrow indicates the spatio-temporal interface moving at $V_{\textrm{interface}}$ here of approximately 8.5 m/s, with blue arrows highlighting the medium’s stretching and drift relative to the laboratory frame. C- Map of the stretch ratio $\lambda$ extracted from the experiment B. D- Repeating the experiment for different $V_{0}$, a linear relation gives $V_{\textrm{interface}}=8.2$ m/s. E- Map of the particle velocity extracted from the experiment B.
@@
$EI\frac{\partial^{4}u_{3}}{\partial{x_{1}}^{4}}-A\sigma\frac{\partial^{2}u_{3}}{\partial{x_{1}}^{2}}+A\rho\frac{\partial^{2}u_{3}}{\partial t^{2}}=0,$ (1)
with $u_{3}$ denoting the out-of plane displacement of the beam, $E$ the Young modulus of the material, $A=wh$ the beam’s cross-section, $I=wh^{3}/12$ the second moment of area, $\sigma$ the applied longitudinal tensile stress and $\rho$ the mass density of the material. This equation immediately reveals that $\sigma$ acts as a tuning parameter: as $\sigma$ varies, a competition between the first and second terms explicitly emerges. At low stress, the propagation is governed by the flexural rigidity $EI$. But, as the stress increases the tensile force $A\sigma$ begins to dominate. Alternatively to the stress, this criterion can also be written in terms of the stretch ratio $\lambda$, invoking Hooke’s law $\sigma=E(\lambda-1)$. Note that this equation remains a simplified model compared to the full description of a true highly stretched beam [34]. Notably, the hyperelastic behaviour as well as the rheology of the materials at play need to be considered for more quantitative measurements. Nevertheless, all these aspects could be incorporated in a Young modulus that would depend on both the frequency and the elongation ratio [34].
@@
Figure 2: Crossing a space-time interface — Space-time diagram (and reciprocal representation) of a flexural wave packet crossing a space-time interface propagating in the opposite direction (A) or in the same direction (B). The strip is pulled at 4 m/s in (A) and 3.3 m/s in (B). The non-invariance of both wavelength (or wavenumber) and period (or frequency) for the transmitted waves is demonstrated. Below are represented the spatio-temporal Fourier transforms of the same data. Superimposed are shown the theoretical dispersion curves before (blue) and after the interface (green).
@@
Figure 1.C shows the strip’s elongation map as a function of space and time. It demonstrates a spatiotemporal transition from initial rest to a stretched configuration. Because $\lambda$ explicitly depends on $V_{0}$, the contrast in elongation between the two regions can be controlled by tuning the motor velocity $V_{0}$. And, as discussed earlier, the amount of elongation directly sets the velocity of the in-plane flexural waves. In comparison to the theoretical model, the interface is not as sharp as a step function but rather exhibits a slower variation from 1 to 1.16. It can be attributed to the acceleration of the motor and to the rheology of the material [39, 40, 34]. A systematic extraction of the stretch ratio for various values of $V_{0}$ is presented in Figure 1.D. Not only it confirms the possibility to create a wide variety of elongation ratios, but the agreement with equation (4) allows to retrieve a precise value of $V_{\textrm{interface}}=8.2$ m/s. This result is in good agreement with the expression of the bar velocity (see above), with $E=67$ kPa and $\rho=1000$ kg/m3, as measured in a previous work [34].
A second crucial aspect to this experimental setup involves a peculiarity of the stretched medium. According to equation (3), after the interface passage, the medium experiences a uniform motion at velocity $V_{0}$. Consequently, not only the medium is stretched but it also undergoes a uniform translation. This appears readily on Figure 1.B, where red dots and blue arrows indicate linear translation to the right. The experimental drifting velocity obtained from the experiment is color-coded in Figure 1.E. After the interface passage, a homogeneous medium drifting at the constant velocity $V_{0}$ is observed in the upper right corner of the figure (once again the interface is not as sharp as a step function and a smooth transition is observed in the figure). This translation significantly affects the propagation of the transverse wave when observed in the laboratory frame. Specifically, due to the translation of the wave’s support, a Doppler shift needs to be accounted for:
@@
These results exemplify the paradigm shift introduced by considering a moving interface. Generally, the invariance of frequency (or equivalently, the temporal period) when crossing a spatial interface is justified by invoking the invariance by time translation along with the linearity of the medium. However, recent experiments involving a temporal interface [44, 45], echoing previous theoretical developments [46, 47], have revealed that the presence of a reflected time-reversed wave can be attributed to the conservation of the wavenumber (or wavelength). In our case, none of these quantities is preserved. It underscores the fundamental changes brought about by the moving interface [30]. This phenomenon can be summarized by the fact that the change of wavenumber is linked to the change of pulsation with respect to the velocity of the interface. From our experimental observations, we evidence the following generalized conservation law:
@@
where $\omega$ corresponds to the angular frequency and $k$ corresponds to the wavenumber of the wave, and is negative if the wave propagates in the negative direction. Note that this expression yields the usual conservation laws regarding space or time interfaces: $V_{\textrm{interface}}=0$ yields $\omega_{2}=\omega_{1}$ (space interface) and $V_{\textrm{interface}}=\infty$ yields $k_{2}=k_{1}$ (time interface).
Despite the broken time and space invariance in this system, the 2D Fourier Transform still yields valuable information about the spectral contents. The results are depicted below the spatio-temporal maps of the waves in Figure 2. For each scenario, two peaks are observed in the reciprocal plane. For the wave propagating in the negative direction, the wavenumbers naturally correspond to negative values. For the strip at rest (blue) as well as for the stretched and moving one (green), all the peaks coincide with the full 3D calculation of the dispersion relations [34, 48]. These calculations fully incorporate the acoustoelastic effect and the non-Hookean nature of the strip. They are more accurate than the oversimplification made in equation (1) and permit to describe accurately the dispersion relation of Medium 2. The transition in the dispersion diagram from the curve of Medium 1 toward the curve of Medium 2 follows the line of slope $-V_{\textrm{interface}}$ and confirms the relation between the wavenumber and frequency shifts of equation (6).
Eventually, the experiment is repeated for various pulling velocities and wavepacket’s frequencies. For each experiment 2D space-time Fourier transforms are applied and the relevant spots of the Fourier transforms are kept. All the results are accumulated and are summarized on the same axes in Figure 3. The initial wavepackets are all aligned with the dispersion relation of flexural waves in a strip almost at rest (black line). A slight difference between the negative and positive wavenumbers is visible because of the gravity that exerts a smaller elongation of the strip at its bottom compared to the top. After crossing the interface, all wavepackets show a change in frequency and wavenumbers. The transmitted wave ends on the dispersion curve of the Medium 2 which depends on the pulling velocity $V_{0}$ with a good agreement with the theoretical predictions (lines on the figure). The transition between the incident wave packet and the transmitted one follows the equation (6) as indicated by the red arrows that are all parallel with the same slope of $-V_{\textrm{interface}}$.
@@
This simple macroscopic experiment effectively demonstrates the key feature of a spatio-temporal interface: the non-conservation of wavenumber and frequency. However, some comments regarding this platform are needed.
@@
Secondly, in this experiment the wave is convected by the motion of the support, resulting in a substantial frequency shift observed in the scenario of figure 2A where the wavepacket travels in the opposite direction compared to the interface. This property can be found in other wave phenomena, such as gravity-capillary waves at the surface of moving fluids [50, 51], acoustic waves in flowing pipes [41, 43], or elastic waves in waveguides coupled to flowing fluids like arteries [52]. These examples could also be promising candidates for investigating static boundaries in the laboratory frame, which would be in translation in the moving frame.
@@
This Letter presents an experiment where a wave packet crosses a moving interface separating two media with distinct propagation properties. This experiment relies on the propagation of a transverse wave in an elastic strip which is dynamically stretched. The observed transmitted wave in our setup exhibits both a frequency shift and a wavenumber shift, driven by the velocity of the interface and the propagation characteristics of the two media.
@@
Acknowledgments —
This work has received support under the program “Investissements d’Avenir” launched by the French Government and partially by the Simons Foundation/Collaboration on Symmetry-Driven Extreme Wave Phenomena. A. D. acknowledges funding from French Direction Générale de l’Armement.